OpenStudy (anonymous):

Solve 6(r-4)=2(r-8)+3r with steps?

OpenStudy (anonymous):

6(r-4)=2(r-8) +3r which is 6r-24 =2r -16 +3r giving r =8...

OpenStudy (anonymous):

thank you